OpenStudy (anonymous):

Simplify 4x/9 + x - 10/9 PLease check my answer! XD

OpenStudy (anonymous):

I got 5x-10/9 is that right?

OpenStudy (anonymous):

no, you get 1 4/9x-10/9

OpenStudy (anonymous):

or 13/9 x -10/9

OpenStudy (anonymous):

its supposed to be adding and subtracting rational expressions

OpenStudy (anonymous):

yes, but you left out the 4x/9 ... you used just 4x... I think

OpenStudy (anonymous):

aww :( i did 4x/9 + x-10/9 = 4x+x-10/9 = 5x-10/9

OpenStudy (anonymous):

4x/9 + x-10/9 = 4/9x+9/9x-10/9 = 13/9x-10/9

OpenStudy (anonymous):

ohhh okay XD thank you!